how much alcohol must a nurse add to 10ml of a 60% alcohol solutions to strengthen it to 90% solution?
Answer by stanbon(75887)

how much alcohol must a nurse add to 10ml of a 60% alcohol solutions to strengthen it to 90% solution?
----------
Equation:
alcohol + alcohol = alcohol
0.60*10 + 1*x = 0.9(10+x)
Multiply thru by 10 to get:
6*10 + 10x = 9*10+9x
x = 3*10
x = 30ml (amt. of pure alcohol needed)
===============
Cheers,
Stan H.
==========


--------------------------------------------------------------------------------